OpenStudy (anonymous):

A rare bacterial culture is being grown in a lab. As the days progress, the cells multiply and grow. After 1 day, there is only 1 cell. After 2 days, there are 5 cells. After 3 days, there are 10 cells. After 4 days, there are 16 cells. Which recursive equation represents the pattern?

OpenStudy (ranga):

The differences form an arithmetic sequence.

OpenStudy (anonymous):

But the differences are different

OpenStudy (ranga):

The differences form the sequence: 4, 5, 6, ...... This is an AP (arithmetic progression). The nth term is given by the formula: a + (n-1)d = 4 + (n-1)*1 = n + 3.

OpenStudy (ranga):

The differences form an AP with the nth difference being (n+3) Put n = 1, 2, 3, 4, etc. and you will get the sequence of differences. On the left hand side of the equation we have a2 - a1 as the first difference. a3 - a2 as the second difference. So the nth difference would be:\[\Large a _{n+1} - a _{n} = n + 3\]

OpenStudy (anonymous):

I'm not understanding any of this tbh

OpenStudy (ranga):

a(subscript n) is the number of bacteria on the nth day a(subscript n+1) is the number of bacteria on the (n+1)th day We start with a1 = 1 And we have found a recursive formula to find a2, a3, a4, etc. Put n = 1, 2, 3, 4 in the equation I have above and you will see what it represents. Put n = 1 a2 - a1= 1 + 3 = 4 a2 = 4 + a1 = 4 + 1 = 5 Put n =2 a3 - a2 = 2 + 3 = 5 a3 = a2 + 5 = 5 + 5 = 10

OpenStudy (anonymous):

but these are set up differently a. an = an - 1 + 2n b. an = an - 1 + n2 c. an = an - 1 + (n + 2) d. an = an - 1 + 2(n + 1)

OpenStudy (ranga):

In my case, n runs from 1 to infinity. In their case, it runs from 2 to infinity. All you have to do is replace n by n-1 in my formula everywhere in my formula to get the answer. They are the same recursive equations but they differ in where n starts.

OpenStudy (anonymous):

would it be an = ((an - 1)) + (n + 2)?

OpenStudy (ranga):

\[\Large a _{n+1} - a _{n} = n + 3 \quad \text{ for n = 1, 2, 3, ...}\]Replace n by (n-1)\[\Large a _{n-1+1} - a _{n-1} = (n-1) + 3\]\[\Large a _{n} - a _{n-1} = n + 2 \quad \text{ for n = 2, 3, 4, ...}\]

OpenStudy (ranga):

Yes.

OpenStudy (ranga):

If I had seen the answer choices along with the problem initially I would have derived the formula with a(n) and a(n-1) from the start. But it is no big deal. Both represent the same but they just differ in the starting value of n.

OpenStudy (anonymous):

Okay thank you!

OpenStudy (ranga):

alright. yw.